# Simulating Non-Markovian Open Quantum Dynamics with Neural Quantum States

### v2

@@ Abstract
Figure 4: Heatmap of $\Delta\bm{K}=\lvert\bm{K}(t_{\rm lo})-\bm{K}(t_{\rm sh})\rvert$ for the open quantum dynamics shown in Fig. 3 at $t_{{\rm sh}}=1.25/\Gamma$ and $t_{{\rm lo}}=5/\Gamma$: (a) $k_{\rm B}T=3.0\,\Gamma$ and (b) $k_{\rm B}T=0.3\,\Gamma$. The vertical axis represents dissipaton decay rates (sorted slow to fast), while the horizontal axis indexes auxiliary nodes. Temporal evolution of auxiliary-layer-averaged weight per dissipaton level, $\Delta K_{j}=\lvert K_{j}(t)-K_{j}(t_{\rm sh})\rvert$: (c) $k_{\rm B}T=3.0\,\Gamma$ and (d) $k_{\rm B}T=0.3\,\Gamma$. The slowest-decaying level is highlighted in red. $\gamma_{j}\equiv{\rm Re}(\gamma^{\sigma}_{j})$ denotes the decay rate of the $j$-th dissipaton level (in unit of $1/\Gamma$). Identical $N_{\rm E}$ is adopted to enable a direct comparison between different temperatures.
As depicted in Fig. 3(d), increasing the number of hidden neurons substantially reduces the loss function $\Delta s^{2}$ over time, indicating that expanding the neural network enhances its representability for quantum correlations. Moreover, Fig. 3(e) demonstrates that the NQS-DQME approach requires significantly fewer dynamical variables than the conventional HEOM method to characterize non-Markovian open quantum dynamics while achieving comparable accuracy.
It is difficult to directly identify traces of non-Markovian memory effects in dynamical variables from HEOM or DQME methods due to their inherent high-dimensional structures. In contrast, the NQS approach recasts these dynamical variables as synaptic weights connecting pairs of neurons, facilitating the extraction of essential non-Markovian features. This unique advantage is exemplified in Fig. 4, which visualizes changes ($\Delta\bm{K}$) in the RBM’s weight matrix $\bm{K}$ between visible nodes (representing dissipaton levels) and auxiliary nodes, from $t_{\rm sh}$ to $t_{\rm lo}$, for the open quantum dynamics shown in Fig. 3. Comparing the $\Delta\bm{K}$ heatmaps in Fig. 4(a) and (b) reveals the outstanding activity of the slowest-decaying dissipaton during long-time dissipative dynamics at low temperature. This observation is further verified by the evolution of the auxiliary-layer-averaged weight for each dissipaton level ($\Delta K_{j}$) in Fig. 4(c) and (d), where the slowest-decaying dissipaton (red line) becomes increasingly prominent as temperature decreases.
@@
Figure 5: (a) Schematic of quantum dissipative dynamics in an OQS comprising two impurities coupled to a reservoir. Spin-exchange interaction (wavy line) and Kondo clouds (gray regions) are shown, where cloud extents indicate Kondo correlation strengths. (b) Temporal evolution of $\dot{S}_{12}$ and (c) $E_{\rm hyb}$ at various temperatures. NQS-DQME results (solid lines, vertically offset for clarity) are benchmarked against HEOM reference values. Inset in (c): $\Delta E_{\rm hyb}$ versus inverse temperature. System parameters (in units of $\Gamma$): $\epsilon_{0}=-U_{0}/2=-6$, and $J=8$; see the SM for details [69].
@@
Case 2: non-Markovian spin relaxation with intertwined Kondo correlations. We now examine a more complex OQS comprising two impurities coupled to a reservoir. This setup directly models recent experimental studies of coherent spin dynamics in surface molecules [14, 15]. The system Hamiltonian is $H_{{}_{\rm S}}(t)=\sum_{\nu=1,2}[\epsilon_{0}(\hat{n}_{\nu\uparrow}+\hat{n}_{\nu\downarrow})+U_{0}\hat{n}_{\nu\uparrow}\hat{n}_{\nu\downarrow}]-\Theta(t-t_{0})J\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{2},$ where the final term activates a ferromagnetic exchange interaction ($J>0$) between the impurities at time $t_{0}$.
@@
Figure 5(a) shows the relaxation dynamics following the sudden activation of the exchange interaction. The initially independent localized spins on the two impurities gradually evolve towards parallel alignment. This results in a localized high-spin state that cannot be fully screened by the reservoir electrons. This underscreening progressively suppresses Kondo correlations during the temporal evolution [79, 80, 7].
@@
Figure 5(b) and (c) depict the time derivative of the inter-impurity spin correlation, $\dot{S}_{12}=\frac{\partial}{\partial t}\langle\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{2}\rangle$, and the impurity-reservoir hybridization energy, $E_{\rm hyb}=\langle H_{{}_{\rm SE}}\rangle$, respectively, for various temperatures. At low temperatures, both quantities exhibit prominent oscillatory lineshapes due to non-Markovian memory effects. We show $\dot{S}_{12}$ instead of ${S}_{12}$ because the oscillatory behavior is more pronounced in the time derivative. The overall change in hybridization energy, $\Delta E_{\rm hyb}=E_{\rm hyb}(t_{\rm lo})-E_{\rm hyb}(t_{0})$, remains positive and increases as temperature decreases (see inset of Fig. 5(c)), indicating weakened Kondo correlations in the long-time limit.
@@
As demonstrated in Fig. 5(b) and (c), the NQS-DQME approach accurately reproduces the reference values. The maximum errors $\mathcal{E}_{{S}_{12}}$ and $\mathcal{E}_{{E}_{\rm hyb}}$ are both below $1\%$ for $t_{\rm lo}>3/\Gamma$. This quantitative agreement is remarkable given that the relaxation dynamics of localized spins are simultaneously influenced by inter-spin correlations and impurity-reservoir Kondo correlations, making them highly intricate. Similar to Case 1, the number of dynamical variables required to capture non-Markovian memory effects is significantly reduced by the NQS approach [69].
Summary. We successfully integrate the NQS approach with the DQME theory, using the concept of dissipatons (statistical quasiparticles encoding environmental memory), to address the challenge of simulating non-Markovian OQSs. The resulting NQS-DQME framework demonstrates remarkable capabilities in capturing the intricate interplay between non-Markovian memory and many-body correlations while enhancing both scalability and interpretability.
While promising, the current NQS-DQME framework requires further development to tackle more complex OQSs. It faces a common bottleneck of NQS approaches: minimization of the loss function $\Delta s^{2}$ involves computationally intensive optimization [56]. Nevertheless, by leveraging advanced algorithms and more sophisticated neural networks, the approach holds great promise for simulating non-Markovian open quantum dynamics with broad-ranging applications.

### v3

@@ Abstract
Figure 4: Visualization of the RBM parameters for the open quantum dynamics shown in Fig. 3. (a,b) Heatmaps of $\lvert\bm{K}(t_{\rm lo})-\bm{K}(t_{\rm sh})\rvert$ at (a) $k_{\rm B}T=3.0\,\Gamma$ and (b) $k_{\rm B}T=0.3\,\Gamma$, for $t_{{\rm sh}}=1.25/\Gamma$ and $t_{{\rm lo}}=5/\Gamma$. The vertical axis indexes dissipaton levels (with decay rates sorted from slowest to fastest), and the horizontal axis indexes the auxiliary nodes. (c,d) Time evolution of the auxiliary-layer-averaged weight difference for the $j$th dissipaton level, $\Delta K_{j}=\lvert K_{j}(t)-K_{j}(t_{\rm sh})\rvert$, at (c) $k_{\rm B}T=3.0\,\Gamma$ and (d) $k_{\rm B}T=0.3\,\Gamma$. Here, $\gamma_{j}\equiv{\rm Re}(\gamma^{\sigma}_{j})$ denotes the decay rate in unit of $1/\Gamma$. The red line tracks the slowest-decaying dissipaton level.
To explore the tradeoff between numerical cost and accuracy, we examine the low temperature regime ($k_{\rm B}T=0.3\,\Gamma$), where non-Markovian effects are prominent. The dissipaton-embedded system comprises $2$ system levels coupled inhomogeneously to $64$ dissipaton levels. Figure 3(c) shows the evolution of the precision metric $\Delta s^{2}$ for different numbers of hidden nodes $N_{\rm h}$. The magnitude of $\Delta s^{2}$ decreases by orders of magnitude with a linear increase in $N_{\rm h}$, demonstrating rapid convergence. Similarly, Fig. 3(d) shows that that accurate results are achieved with relatively few Monte Carlo samples ($N_{\rm MC}$). Furthermore, Fig. 3(e) demonstrates that the NQS-DQME approach achieves high accuracy ($\Delta s^{2}<10^{-8}$) while utilizing drastically fewer dynamical variables than the conventional HEOM method. This comparison is conducted under equivalent conditions, as the same symmetry and sparsity relations are incorporated into both methods. These results collectively highlight the outstanding expressive power of the RBM for characterizing non-Markovian open quantum dynamics.
An important hypothesis in open quantum dynamics is that non-Markovian memory is crucial for long-time phenomena like the formation of Kondo-correlated states. However, directly verifying this by tracking dynamical variables in conventional methods such as the HEOM or the original DQME is prohibitively difficult due to their high dimensionality. In contrast, the NQS-DQME framework makes such tracking feasible, since the RDT is represented by weight matrices that form a much lower-dimensional parameter space.
This capability is demonstrated in Fig. 4, which visualizes the evolution of the RBM weight matrix $\bm{K}$ connecting dissipaton-level nodes to auxiliary nodes for the open quantum dynamics shown in Fig. 3. A comparison of the heatmaps in Fig. 4(a) and (b) shows that the weight changes, $\lvert\bm{K}(t_{\rm lo})-\bm{K}(t_{\rm sh})\rvert$, are significantly more pronounced for the slowest-decaying dissipaton level at the lower temperature. This is indicated by the darker top row in Fig. 4(b). This observation is further confirmed by the time evolution of the auxiliary-layer-averaged weight difference $\Delta K_{j}(t)$ in Fig. 4(c) and (d), where the slowest-decaying dissipaton level (red line) becomes increasingly prominent as temperature decreases, directly linking the non-Markovian memory to the emerging Kondo effect.
@@
Figure 5: (a) Schematic of quantum dissipative dynamics for two impurities coupled to a reservoir. Spin-exchange interaction (wavy line) and Kondo clouds (gray regions) are shown, where cloud extents indicate Kondo correlation strengths. Time evolution of (b) $\dot{S}_{12}$ and (c) $E_{\rm hyb}$ at various temperatures. NQS-DQME results (solid lines, vertically offset for clarity) are benchmarked against HEOM reference values. Inset in (c): $\Delta E_{\rm hyb}=E_{\rm hyb}(t_{\rm lo})-E_{\rm hyb}(t_{0})$ versus inverse temperature. The increase in $\Delta E_{\rm hyb}$ with decreasing temperature indicates weakened Kondo correlations in the long-time limit. System parameters (in units of $\Gamma$): $\epsilon_{0}=-U_{0}/2=-6$, and $J=8$; see the SM for details [SM].
@@
Case 2: non-Markovian spin relaxation with intertwined Kondo correlations. We now examine a more complex OQS comprising two impurities coupled to a reservoir. This setup directly models recent experimental studies of coherent spin dynamics in surface molecules [doi:10.1126/science.aay6779, doi:10.1126/science.abg8223]. The system Hamiltonian is $H_{{}_{\rm S}}(t)=\sum_{\nu=1,2}[\epsilon_{0}(\hat{n}_{\nu\uparrow}+\hat{n}_{\nu\downarrow})+U_{0}\hat{n}_{\nu\uparrow}\hat{n}_{\nu\downarrow}]-\Theta(t-t_{0})J\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{2},$ where the final term activates a ferromagnetic exchange interaction ($J>0$) between the impurities at time $t_{0}$.
@@
Figure 5(a) shows the relaxation dynamics following the sudden activation of the exchange interaction. The initially independent localized spins on the two impurities gradually evolve towards parallel alignment. This results in a localized high-spin state that cannot be fully screened by the reservoir electrons. This underscreening progressively suppresses Kondo correlations during the temporal evolution [PhysRevLett.103.197202, parks2010mechanical, li2020molecular].
@@
Figure 5(b) and (c) depict the time derivative of the inter-impurity spin correlation, $\dot{S}_{12}=\frac{\partial}{\partial t}\langle\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{2}\rangle$, and the impurity-reservoir hybridization energy, $E_{\rm hyb}=\langle H_{{}_{\rm SE}}\rangle$, respectively, at various temperatures. We plot $\dot{S}_{12}$ rather than ${S}_{12}$ to highlight its more pronounced oscillatory behavior. At low temperatures, both quantities exhibit strong oscillations, a hallmark of non-Markovian memory effects.
@@
The NQS-DQME results are in excellent agreement with the reference values, with maximum errors $\mathcal{E}_{{S}_{12}}$ and $\mathcal{E}_{{E}_{\rm hyb}}$ both below $1\%$ for $t_{\rm lo}>3/\Gamma$. This quantitative agreement is remarkable, as the spin relaxation dynamics are governed by the complex interplay of direct inter-impurity correlations and impurity-reservoir Kondo correlations. Furthermore, similar to Case 1, the NQS representation achieves this accuracy with a significantly reduced number of dynamical variables compared to conventional approaches [SM].
Summary. We have developed a NQS framework for the efficient simulation of non-Markovian open quantum dynamics by integrating the dissipaton concept into the ansatz. This NQS-DQME approach provides a compact and physically interpretable epresentation of the RDT, enabling the simultaneous treatment of complex environmental memory and many-body correlations with significantly enhanced computational efficiency.
While the current framework shares a common challenge with many NQS methods, i.e., the computationally demanding minimization of the loss function $\Delta s^{2}$ [PhysRevLett.122.250503]. This bottleneck can be addressed using advanced optimizers and more sophisticated network architectures. Given its proven scalability and interpretability, the NQS-DQME approach presents a promising and versatile platform for exploring rich quantum phenomena in correlated, non-Markovian settings.